# On Geometry and Topology of Hessian Manifolds

### v3

@@ 2.3.2 The canonical line bundle
We now introduce the notion of Koszul forms of a Hessian line bundle.
@@
$\varkappa|_{U_{\alpha}}:=-\frac{1}{2}d\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
@@
for every vector field $X$ on $M$, where $|h|$ is the determinant of the matrix of which $(i,j)$-entry is $h(\partial/\partial x^{i},\partial/\partial x^{j})$ for some oriented affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$.
@@
In [12], H.Shima and K.Yagi also established the following criterion for flatness of Hessian manifolds in terms of the first Koszul form.
@@
Since $\mathrm{H}^{1}(M;\mathbb{R})=0$, by Theorem2.27 we have that the Picard group of $M$ is isomorphic to $\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})$, and hence is 2-torsion. In particular, the canonical line bundle $K$ of $(M,\nabla)$ satisfies $K^{\otimes 2}=0$. Since $M$ is orientable, the $\mathbb{R}$-line bundle $\bigwedge^{n}(T^{*}M)$ is trivial. Therefore, by virtue of Lemma2.39, the canonical line bundle of $M$ is trivial as a flat $\mathbb{R}$-line bundle. Corollary2.32 then implies that $M$ admits a volume form $\Omega$ such that $\nabla\Omega=0$. The desired result follows immediately from Theorem2.38. ∎
@@
Let $((M,\nabla),h)$ be an orientable compact Hessian manifold. Then either $(M,h)$ is a flat Riemannian manifold, or there exists a fiber bundle $\varphi\colon M\rightarrow\mathbb{S}^{1}$ with connected fiber.
@@
of homotopy groups yields that $F$ is also a connected topological space. ∎
Corollary 2.44.
Let $((M,\nabla),h)$ be an closed orientable Hessian $3$-manifold. Then either $(M,h)$ is a flat Riemannian manifold, or $M$ is a $\Sigma_{g}$-bundle over $\mathbb{S}^{1}$, where $\Sigma_{g}$ is the closed orientable surface of genus $g\geq 1$.
Proof.
Assume $(M,h)$ is not flat. Then by Theorem2.43, $M$ is a $\Sigma$-bundle over $\mathbb{S}^{1}$ for some closed surface $\Sigma$. Since $\pi_{1}(M)$ is torsion-free by Proposition2.19 and $\pi_{1}(\Sigma)$ is a subgroup of $\pi_{1}(M)$ by the long exact sequence
$\cdots\longrightarrow\pi_{2}(\mathbb{S}^{1})\longrightarrow\pi_{1}(\Sigma)\longrightarrow\pi_{1}(M)\longrightarrow\cdots$
of homotopy groups, we have that $\pi_{1}(\Sigma)$ is torsion-free, and hence $\Sigma$ is orientable. Since $M$ is aspherical by Corollary2.14, we obtain that $\Sigma$ ditto. Therefore, in particular, $\Sigma$ is not homeomorphic to $\mathbb{S}^{2}$. This concludes the proof. ∎
2.3.3 The obstruction line bundle
Resembling the notion of positive holomorphic line bundles in complex geometry, we now define and study the positivity of a Hessian line bundle.
@@ 2.3.3 The obstruction line bundle
By the very definition, if an affine manifold $(M,\nabla)$ admits a positive flat line bundle $(\mathscr{L},D)$, then $((M,\nabla),\nabla\varkappa)$ is a Hessian manifold of Koszul type, where $\varkappa$ is the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ and $\langle\cdot,\cdot\rangle$ is a bundle metric of $\mathscr{L}\rightarrow M$ such that $\nabla\varkappa>0$.
Besides the canonical line bundle, another important flat $\mathbb{R}$-line bundle over a Hessian manifold is its obstruction line bundle, which we shall now define.
Definition 2.46.
Let $((M,\nabla),h)$ be a Hessian manifold. Consider an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a smooth function $f_{\alpha}\in C^{\infty}(U_{\alpha})$, such that for each $\alpha\in I$ it holds that $h_{\alpha}=\nabla(df_{\alpha})$, where $h_{\alpha}$ is the restriction of $h$ on $U_{\alpha}$. The flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$, where $\mathscr{L}\rightarrow M$ is the $\mathbb{R}$-line bundle over $M$ defined by the 1-cocycle $\{\operatorname{exp}(f_{\alpha}-f_{\beta})\}_{\alpha,\beta\in I}$ and where $D$ is the flat connection on $\mathscr{L}\rightarrow M$ defined by the connection 1-forms $\{df_{\alpha}\}_{\alpha\in I}$, is termed the obstruction line bundle of $((M,\nabla),h)$.
Attentive readers are invited to check the fact that the obstruction line bundle of a Hessian manifold is uniquely defined up to gauge equivalence.
Also note, for an obstruction line bundle $(\mathscr{L},D)$, since the structure group of $\mathscr{L}$ reduces to $\operatorname{SO}(1)$, we have that $\mathscr{L}$ is always a trivial $\mathbb{R}$-line bundle.
Now, we can reformulate the conditions on the existence of global potential for a Hessian manifold in terms of its obstruction line bundle.
@@
Let $((M,\nabla),h)$ be a Hessian manifold of Koszul type, and $(\mathscr{L},D)$ its obstruction line bundle. Then there exists a bundle metric $\langle\cdot,\cdot\rangle$ on $\mathscr{L}\rightarrow M$ such that the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is a $1$-form potential of $((M,\nabla),h)$. Moreover, if it is assumed further that there exists $f\in C^{\infty}(M)$ such that $h=\nabla(df)$, then $(\mathscr{L},D)$ is a trivial flat $\mathbb{R}$-line bundle.
Proof.
Let $\eta$ be a $1$-form potential of $((M,\nabla),h)$. Since $\eta$ is closed by Proposition2.9, by Poincaré’s lemma there exists an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a smooth function $f_{\alpha}\in C^{\infty}(U_{\alpha})$, such that for each $\alpha\in I$ it holds that $\eta_{\alpha}=df_{\alpha}$, where $\eta_{\alpha}$ is the restriction of $\eta$ on $U_{\alpha}$. Then $\mathscr{L}\rightarrow M$ is the $\mathbb{R}$-line bundle over $M$ defined by the 1-cocycle $\{\operatorname{exp}(f_{\alpha}-f_{\beta})\}_{\alpha,\beta\in I}$ and $D$ is the flat connection on $\mathscr{L}\rightarrow M$ defined by the connection 1-forms $\{df_{\alpha}\}_{\alpha\in I}$. For any $\alpha,\beta\in I$, since $df_{\alpha}-df_{\beta}=0$, we have that the function $\operatorname{exp}(f_{\alpha}-f_{\beta})$ is constant.
Let $\{s_{\alpha}\}_{\alpha\in I}$ be an affine local frame of $(\mathscr{L},D)$ and for each $\alpha\in I$ denote by $U_{\alpha}$ the domain of definition of $s_{\alpha}$, so that for each $\alpha,\beta\in I$ there exists a constant $g_{\alpha\beta}\in\mathbb{R}^{\times}$ such that $s_{\beta}(p)=s_{\alpha}(p)g_{\alpha\beta}$ for all $p\in U_{\alpha}\cap U_{\beta}$. Define a bundle metric $\langle\cdot,\cdot\rangle$ on $\mathscr{L}$ by
for every $\alpha\in I$. Then, $\langle\cdot,\cdot\rangle$ is well-defined, and by construction the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is precisely $\eta$.
Now, assume further that $\eta$ is exact. Then $D=d+\eta$ differs from $d$ by an exact 1-form. By definition, $(\mathscr{L},D)$ is a trivial flat $\mathbb{R}$-line bundle. ∎
Corollary 2.48.
Let $((M,\nabla),h)$ be a Hessian manifold. Then the obstruction line bundle of $((M,\nabla),h)$ is a trivial flat $\mathbb{R}$-line bundle if and only if there exists $f\in C^{\infty}(M)$ such that $h=\nabla(df)$.
Proof.
By Proposition2.47, it remains to prove that if obstruction line bundle $(\mathscr{L},D)$ of $((M,\nabla),h)$ is a trivial flat $\mathbb{R}$-line bundle then there exists $f\in C^{\infty}(M)$ such that $h=\nabla(df)$. But, if
$(\mathscr{L},D)=(M\times\mathbb{R},d+df)$
for some $f\in C^{\infty}(M)$, then it is readily seen that $h=\nabla(df)$. ∎
Corollary 2.49.
@@
This follows from Proposition2.47 and the very definition of a positive flat line bundle. ∎
@@ 3.1 Digression on Complex Geometry
for all $i=1,\dots,n$. By the classical construction in [23], the linear operator $J_{\nabla}$ is a well-defined integrable complex structure, i.e. $(TM,J_{\nabla})$ is a complex manifold. The complex manifold $(TM,J_{\nabla})$ is said to be the Sasaki lifting of $(M,\nabla)$.
@@
Let $(M,\nabla)$ be an affine manifold, and $(TM,J_{\nabla})$ its Sasaki lifting. Suppose that $(\mathscr{L},D)$ is a positive flat line bundle over $(M,\nabla)$. Then the holomorphic pullback $(\mathscr{L}_{\mathbb{C}},\bar{\partial}_{D})$ of $(\mathscr{L},D)$ along $\nabla$ is a positive holomorphic line bundle over $(TM,J_{\nabla})$.
@@ 3.4 Para-Kähler Host Theorem
Let $((M,\omega),g)$ be a flat para-Kähler manifold, and let $D$ be the Levi-Civita connection of $(M,g)$. Suppose that $x^{1},\dots,x^{n},y_{1},\dots,y_{n}$ and $\tilde{x}^{1},\dots,\tilde{x}^{n},\tilde{y}_{1},\dots,\tilde{y}_{n}$ are both affine Darboux coordinates of $((M,\omega),g)$ at the vicinity of $p\in M$. Then there exist affine functions $u_{1},\dots u_{n}\colon\mathbb{R}^{n}\rightarrow\mathbb{R}$ such that $\tilde{x}^{i}=u_{i}(x^{1},\dots,x^{n})$ for any $i\in\{1,\dots,n\}$.
@@ 3.5 Hyperbolic Affine Manifolds
Combine Corollary2.49&3.30 with Theorem3.43. ∎
@@
Let $(M,\nabla)$ be an oriented affine manifold, and let
[…]
be the holonomy of $\nabla$. The homomorphism
$\pi_{1}(M)\rightarrow\mathbb{R},\quad g\mapsto\operatorname{log}(\operatorname{det}(\rho(g)))$
as a group cohomology class in $\mathrm{H}^{1}(\pi_{1}(M);\mathbb{R})$, is termed the canonical class of $(M,\nabla)$.
The canonical class of an oriented affine manifold is by no means an unfamiliar object: We now show that, when the oriented affine manifold is equipped with a complete Hessian metric, its first Koszul form represents precisely the canonical class.
Lemma 3.46.
Let $((M,\nabla),h)$ be a complete oriented Hessian manifold. Then, under the identification $\operatorname{H}^{1}(M;\mathbb{R})=\operatorname{H}^{1}(\pi_{1}(M);\mathbb{R})$, the cohomology class represented by the first Koszul form of $((M,\nabla),h)$ is the canonical class of $(M,\nabla)$.
Proof.
Denote by $\varkappa\in\Omega^{1}(M)$ the first Koszul form of $((M,\nabla),h)$. Since
$\nabla_{X}\operatorname{vol}_{h}=\varkappa(X)\operatorname{vol}_{h}$
for every vector field $X$ on $M$, it follows that $\varkappa$ is the connection 1-form of the induced connection $\nabla$ on $\bigwedge^{n}T^{*}M$. Therefore
$\operatorname{log}(\operatorname{det}(\rho(g)))=\int_{\gamma}\varkappa$
[…]
@@
Let $((M,\nabla),h)$ be a complete oriented Hessian manifold. Suppose that the canonical class of $(M,\nabla)$ is a rational cohomology class. Then the canonical class of $(M,\nabla^{*})$ is also rational.
@@
The desired result follows immediately from Lemma3.46 and the fact that the determinant of $h$ in affine coordinates of $(M,\nabla^{*})$ is the reciprocal of the determinant of $h$ in affine coordinates of $(M,\nabla)$. ∎
We shall show that, on an oriented compact manifold, an hyperbolic affine structure can always be deformed into another hyperbolic affine structure of which canonical class is rational.
@@
3.6 Riemannian Geometric Aspect

### v4

@@ 2.3.2 Applications of Koszul Forms
In this section, we shall explore various applications of flat $\mathbb{R}$-line bundles and their first Koszul forms, including a powerful structure theorem of compact Hessian manifolds.
We now introduce the notion of the first Koszul form of a Hessian line bundle.
@@
$\varkappa|_{U_{\alpha}}:=\frac{1}{2}d\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
@@
A natural question is to compare the first Koszul class of a flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ and the cohomology class of the first Koszul form of Hessian line bundle $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$, where $\langle\cdot,\cdot\rangle$ is some bundle metric on $\mathscr{L}$. Surprisingly, the cohomology class of the first Koszul form is independent of the auxiliary bundle metric and always coincides with the first Koszul class.
Proposition 2.40.
Let $M$ be a differentiable manifold, and let $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ be a Hessian line bundle over $M$. Let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$, then we have $\operatorname{sh}(D)=[\varkappa]\in\mathrm{H}^{1}(M;\mathbb{R})$ is the first Koszul class of $(\mathscr{L},D)$.
Proof.
Let $\{s_{\alpha}\}_{\alpha\in I}$ be an affine local frame field for $(\mathscr{L},D)$ and let $\{g_{\alpha\beta}\}_{\alpha,\beta\in I}$ be its gluing 1-cocycle.
By the very definition, the first Koszul class $\operatorname{sh}(D)$ sends the class of the cocycle $\{g_{\alpha\beta}\}_{\alpha,\beta\in I}$ in $\mathrm{H}^{1}(M;\mathbb{R}^{\times})$ to the class of the cocycle $\{\operatorname{log}|g_{\alpha\beta}|\}_{\alpha,\beta\in I}$ in $\mathrm{H}^{1}(M;\mathbb{R})$.
By Definition 2.39, the first Koszul form $\varkappa$ is a global 1-form defined locally by
$\frac{1}{2}d\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
for $\alpha\in I$. The de Rham cohomology class $\varkappa]$ corresponds to a Čech cohomology class in $\mathrm{H}^{1}(M;\underline{\mathbb{R}})$ via the de Rham isomorphism. For any $\alpha\in I$, define
$f_{\alpha}:=\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle).$
Then for any $\alpha\in I$, we have
$d(f_{\alpha})=\frac{1}{2}d\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle).$
Straightforward computation yields that
$f_{\beta}-f_{\alpha}$ $=\frac{1}{2}\operatorname{log}(\langle s_{\beta},s_{\beta}\rangle)-\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
$=\frac{1}{2}\operatorname{log}(\langle s_{\alpha}g_{\alpha\beta},s_{\alpha}g_{\alpha\beta}\rangle)-\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
$=\frac{1}{2}\operatorname{log}(g_{\alpha\beta}^{2}\langle s_{\alpha},s_{\alpha}\rangle)-\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
$=\frac{1}{2}\left(\operatorname{log}(g_{\alpha\beta}^{2})+\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)\right)-\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
$=\frac{1}{2}\operatorname{log}(g_{\alpha\beta}^{2})$
$=\operatorname{log}|g_{\alpha\beta}|.$
Therefore, a Čech cocycle representing $\varkappa]$ is $\{\operatorname{log}|g_{\alpha\beta}|\}_{\alpha,\beta\in I}$, and hence $\operatorname{sh}(D)=[\varkappa]$. ∎
On a Hessian manifold, we have a canonical Koszul form, which we will now define.
@@
for every vector field $X$ on $M$, where $|h|$ is the determinant of the matrix of which $(i,j)$-entry is $h(\partial/\partial x^{i},\partial/\partial x^{j})$ for some oriented affine coordinates $x^{1},\dots,x^{n}$ of the oriented affine manifold $(M,\nabla)$.
@@
In [12], H.Shima and K.Yagi established the following criterion for flatness of Hessian manifolds in terms of the first Koszul form.
@@
Since $\mathrm{H}^{1}(M;\mathbb{R})=0$, by Proposition 2.32 we have that the Picard group of $M$ is isomorphic to $\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})$, and hence is 2-torsion. In particular, the canonical line bundle $K$ of $(M,\nabla)$ satisfies $K^{\otimes 2}=0$. Since $M$ is orientable, the $\mathbb{R}$-line bundle $\bigwedge^{n}(T^{*}M)$ is trivial. Therefore, by virtue of Lemma 2.35, the canonical line bundle of $M$ is trivial as a flat $\mathbb{R}$-line bundle. Corollary 2.37 then implies that $M$ admits a volume form $\Omega$ such that $\nabla\Omega=0$. The desired result follows immediately from Theorem 2.44. ∎
@@
Let $((M,\nabla),h)$ be an oriented compact Hessian manifold. Then either $(M,h)$ is a flat Riemannian manifold, or there exists a fiber bundle $\varphi\colon M\rightarrow\mathbb{S}^{1}$ with connected and oriented fiber.
@@
of homotopy groups yields that $F$ is also a connected topological space.
Moreover, since $\varphi\colon M\rightarrow\mathbb{S}^{1}$ is a submersion between compact differentiable manifolds, we have that $F$ is also a compact differentiable manifold.
By the very definition of $\varphi$, it is readily seen that the vector field $\hat{\varkappa}^{\sharp}$ is transverse to the fibers of $\varphi\colon M\rightarrow\mathbb{S}^{1}$. Consider the (n1)(n-1)-form $\Omega$ on $M$ defined by
$\Omega(X_{1},\dots,X_{n-1}):=\operatorname{vol}_{h}(X_{1},\dots,X_{n-1},\hat{\varkappa}^{\sharp})$
where $X_{1},\dots,X_{n-1}$ are arbitrary vector fields on $M$. Then by construction $\Omega|_{F}$ is nowhere vanishing. Therefore we conclude that $F$ is oriented. ∎
As yet another application of the first Koszul forms, resembling the notion of positive holomorphic line bundles in complex geometry, we now define the positivity of a flat $\mathbb{R}$-line bundle, and use it to characterize Hessian metrics of Koszul type.
@@
Let $(M,\nabla)$ be an arbitrary but fixed affine manifold.
Assume that there exists a Hessian metric $h$ on $(M,\nabla)$ and a closed $1$-form $\eta\in\Omega^{1}(M)$ such that $h=\nabla\eta$. Consider an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a smooth function $f_{\alpha}\in C^{\infty}(U_{\alpha})$, such that for each $\alpha\in I$ it holds that $\eta_{\alpha}=df_{\alpha}$, where $\eta_{\alpha}$ is the restriction of $\eta$ on $U_{\alpha}$. The flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$,
Let $\mathscr{L}\rightarrow M$ be the $\mathbb{R}$-line bundle over $M$ defined by the 1-cocycle $\{\operatorname{exp}(f_{\alpha}-f_{\beta})\}_{\alpha,\beta\in I}$ and let $D$ be the flat connection on $\mathscr{L}\rightarrow M$ defined by the connection 1-forms $\{df_{\alpha}\}_{\alpha\in I}$.
Let $\{s_{\alpha}\}_{\alpha\in I}$ be an affine local frame of $(\mathscr{L},D)$. Without loss of generality, for each $\alpha\in I$, let $U_{\alpha}$ be the domain of definition of $s_{\alpha}$. Define a bundle metric $\langle\cdot,\cdot\rangle$ on $\mathscr{L}$ by
for every $\alpha\in I$. Then, $\langle\cdot,\cdot\rangle$ is well-defined, and by construction the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is precisely $\eta$. Therefore $\nabla\eta=h>0$, and hence $(\mathscr{L},D)$ is positive.
Conversely, if $(M,\nabla)$ admits a positive flat line bundle $(\mathscr{L},D)$, then by definition there exists a bundle metric $\langle\cdot,\cdot\rangle$ such that $h:=\nabla\varkappa>0$ is a Hessian metric of Koszul type, where $\varkappa$ is the first Koszul form of the Hessian line bundle $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$. ∎
@@ 3.1 Digression on Complex Geometry
for all $i=1,\dots,n$. By the classical construction in [24], the linear operator $J_{\nabla}$ is a well-defined integrable complex structure, i.e. $(TM,J_{\nabla})$ is a complex manifold. The complex manifold $(TM,J_{\nabla})$ is said to be the Sasaki lifting of $(M,\nabla)$, and the holomorphic coordinates
$z^{1}:=x^{1}+\sqrt{-1}v^{1},\dots,z^{n}:=x^{n}+\sqrt{-1}v^{n}$
of $(TM,J_{\nabla})$ are termed the canonical complexifications of the affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$.
@@
Let $(M,\nabla)$ be an affine manifold, and $(TM,J_{\nabla})$ its Sasaki lifting. Suppose that $(\mathscr{L},D)$ is a positive flat line bundle over $(M,\nabla)$. Then the holomorphic pullback $(\mathscr{L}_{\mathbb{C}},\bar{\partial}_{D})$ of $(\mathscr{L},D)$ along $\nabla$ is a positive holomorphic line bundle over the complex manifold $(TM,J_{\nabla})$.
@@
After the notion of pullback of line bundles being successfully defined, we also want to discover a proper way to lift smooth functions from the base. Instead of naively pullback smooth functions on an affine manifold to its tangent bundle via the bundle projection, there exists a much more purposeful method of lifting smooth functions.
Proposition 3.3.
Let $(M,\nabla)$ be an affine manifold, and $(TM,J_{\nabla})$ its Sasaki lifting. Let $f\colon M\rightarrow\mathbb{R}$ and $F\colon TM\rightarrow\mathbb{C}$ be smooth functions satisfying
$F(x^{1},\dots,x^{n};v^{1},\dots,v^{n})=f(x^{1},\dots,x^{n})+\sqrt{-1}\sum_{i=1}^{n}\frac{\partial f}{\partial x^{i}}(x^{1},\dots,x^{n})v^{i}$
where here
$z^{1}=x^{1}+\sqrt{-1}v^{1},\dots,z^{n}=x^{n}+\sqrt{-1}v^{n}$
are the canonical complexifications of some affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$. Then $f\in C^{\infty}(M)$ is affine if and only of $F\in C^{\infty}(TM)\otimes\mathbb{C}$ is holomorphic.
Proof.
Apply Cauchy-Riemann equation to the specific complex structure $J_{\nabla}$. ∎
Proposition 3.3 then motivates the following definition for general smooth functions.
Definition 3.4.
Let $(M,\nabla)$ be an affine manifold, and let $(TM,J_{\nabla})$ be its Sasaki lifting. Let
$z^{1}=x^{1}+\sqrt{-1}v^{1},\dots,z^{n}=x^{n}+\sqrt{-1}v^{n}$
be the canonical complexifications of some affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$. For any smooth function $f\in C^{\infty}(M)$ on $M$, The complex-valued smooth function $F\in C^{\infty}(TM)\otimes\mathbb{C}$ defined by
$F(x^{1},\dots,x^{n};v^{1},\dots,v^{n}):=f(x^{1},\dots,x^{n})+\sqrt{-1}\sum_{i=1}^{n}\frac{\partial f}{\partial x^{i}}(x^{1},\dots,x^{n})v^{i}$
is termed the canonical complexification of $f$ along $\nabla$.
Note that, the notion of canonical complexifications of affine coordinates is consistent with the definition of canonical complexifications of smooth functions, which also justifies the usage of terminology.
@@ 3.4 Para-Kähler Host Theorem
Let $((M,\omega),g)$ be a flat para-Kähler manifold, and let $D$ be the flat Levi-Civita connection of $(M,g)$. Suppose that both $x^{1},\dots,x^{n},y_{1},\dots,y_{n}$ and $\tilde{x}^{1},\dots,\tilde{x}^{n},\tilde{y}_{1},\dots,\tilde{y}_{n}$ are affine Darboux coordinates of $((M,\omega),g)$ at the vicinity of $p\in M$. Then there exist locally defined affine functions $u_{1},\dots u_{n}$ on $(M,D)$ such that $\tilde{x}^{i}=u_{i}(x^{1},\dots,x^{n})$ for $i=1,\dots,n$.
@@ 3.5 Hyperbolic Affine Manifolds
Combine Proposition 2.503.32 with Theorem 3.45. ∎
@@
The first Koszul class of the canonical line bundle of an affine manifold $(M,\nabla)$ is said to be the canonical class of $(M,\nabla)$.
We recall that, by Proposition 2.40, when the affine manifold under consideration is equipped with a Hessian metric, its first Koszul form represents precisely the canonical class.
@@
Let $((M,\nabla),h)$ be a complete oriented Hessian manifold. Suppose that the canonical class of $(M,\nabla)$ is $K$. Then the canonical class of $(M,\nabla^{*})$ is K-K.
@@
The desired result follows immediately from Proposition 2.40 & 2.33 and the fact that the determinant of $h$ in affine coordinates of $(M,\nabla^{*})$ is the reciprocal of the determinant of $h$ in affine coordinates of $(M,\nabla)$. ∎
We shall show that, on an oriented compact manifold, an hyperbolic affine structure can always be perturbed into another hyperbolic affine structure of which canonical class is rational.
@@
4 Geometric Analysis of Hessian Metrics
4.1 Effects on Curvature Tensors